A parallel polarized wave is incident from air into paraffin having relative permittivity 3, the value of its Brewster angle is


1
Expert's answer
2021-01-13T11:45:02-0500

Answer

Brewster angle is given by

"\theta=\tan^{-1} (\sqrt{\frac{\epsilon_{r'}}{ 
 \epsilon_{r} }})"

"=\tan^{-1} (\sqrt{\frac{3}{1}}) =60^°"